OpenStudy (anonymous):

the difference of two numbers is 3 ; and three times the greater exceeds twice the less by 18 . find the numbers

OpenStudy (mokeira):

let the numbers be x and y \[x-y=3\] \[3x=2y+18\] \[x=3+y\] substitute \[3(3+y)=2y+18\] solve for y \[9+3y=2y+18\] \[y=9\] earlier on we found that \[x=3+y\] we can now find x \[x=3+9=12\]

OpenStudy (mokeira):

@keylaunie